A function is given. Determine the average rate of change of the function between the given values of the variable.
f(x) = 3x^2; x = 4, x = 4 + h

Answer by stanbon(75887)

A function is given. Determine the average rate of change of the function between the given values of the variable.
f(x) = 3x^2; x = 4, x = 4 + h
-----
Average = [f(4+h)-f(4)]/[(4+h)-4] = [f(4+h)-f(4)]/h
---------
f(4+h) = 3(4+h)^2 = 3(16+8h+h^2) = 48+24h+3h^2
f(4) = 3*4^2 = 48
====
Average = [(48+24h+3h^2)-48]/h = (24h+3h^2)/h = 24+3h
=======================================================
Cheers,
Stan H.
=============